# NeoVerse: Enhancing 4D World Model with in-the-wild Monocular Videos

### v1

@@ 4.3 Qualitative Evaluation and Analysis
It is noteworthy that our model also has better pose prediction accuracy. In Fig. 5, the compared method [anysplat] shows a field of view (images with red boundaries) inconsistent with the ground truth, which is caused by inaccurate pose prediction.
@@ 4.5 Applications
Since our model has a binary mask condition and a textual condition, it can edit videos with the help of a video segmentation model [sam2], demonstrated in Fig. 10.
@@ 5 Conclusion and Limitations
[…]
@@ Appendix A Implementation Details
The transformer decoders in the bidirectional motion-encoding branch follow the design of DUSt3R [dust3r], where each decoder block consists of a self-attention layer for intra-frame spatial modeling and a cross-attention layer for inter-frame temporal modeling. Finally, two DPT [dpt] heads are employed to predict the forward and backward motions, respectively. Here, we define the forward/backward velocities $\{\boldsymbol{v}_{i}^{+},\boldsymbol{v}_{i}^{-}\}$ as the 3D displacements from the current frame to the next/previous frame in the camera coordinate.
@@
The multiple encoders for multi-modal conditions are implemented with 1) VAE [wan] encoder for RGB images and depth maps, 2) convolutional layers with $8\times$ spatial and $4\times$ temporal compression ratio for masks and plüker embeddings. During the generation training stage, only convolutional layers are trainable while the VAE encoder is frozen.
@@ Appendix B Training Details
To ensure compatibility with the patch size of DINOv2 [dinov2] in the reconstruction model ($\times 14$ downsampling) and the VAE in the generation model ($\times 8$ compression), we resize all input videos to have a longest edge of 560 pixels during reconstruction training, and a fixed resolution of $336\times 560$ during generation training.
@@
PointOdyssey [pointodyssey] 131
DynamicReplica [dynamicstereo] 483
Kubric [kubric] 5.7K
Spring [spring] 37
VKITTI2 [vkitti2] 50
Waymo [waymo] 798
TartanAir [tartanair] 369
BEDLAM [bedlam] 10.4K
MVS-Synth [deepmvs] 120
GFIE [gfie] 81
HOI4D [hoi4d] 3.0K
CoP3D 2.8K
DL3DV [dl3dv] 6.4K
Scannet++ [scannet++] 853
ARKitScenes [arkitscenes] 4.5K
HyperSim [arkitscenes] 457
MapFree [mapfree] 460
SpatialVID [spatialvid] 371.3K
@@ Appendix D Evaluation Protocol
We evaluate static reconstruction performance on VRNeRF [vrnerf] and Scannet++ [scannet++].
@@
For dynamic reconstruction on ADT [adt], we follow 4DGT [4dgt] to evaluate the same 4 scenes:
@@
For DyCheck [dycheck], we evaluate 5 scenes (apple, block, paper-windmill, spin, teddy). We sample 64 consecutive timestamps for each scene, using 32 frames (stride 2) from a casually-captured video (camera 0) for reconstruction and the complete 64 frames from another fixed-camera video (camera 1) for testing.
Appendix E Limitations and Failure Cases
@@
Appendix F Additional Qualitative Results

### v2

@@ 4.3 Qualitative Evaluation and Analysis
It is noteworthy that our model also has better pose prediction accuracy. In Fig. 5, the compared method [32] shows a field of view (images with red boundaries) inconsistent with the ground truth, which is caused by inaccurate pose prediction.
@@ 4.5 Applications
Since our model has a binary mask condition and a textual condition, it can edit videos with the help of a video segmentation model [57], demonstrated in Fig. 10.
@@ 5 Conclusion and Limitations
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (No. 62320106010) and in part by Beijing Natural Science Foundation (No. L257004, No. L257015).
@@ Appendix A Implementation Details
The transformer decoders in the bidirectional motion-encoding branch follow the design of DUSt3R [71], where each decoder block consists of a self-attention layer for intra-frame spatial modeling and a cross-attention layer for inter-frame temporal modeling. Finally, two DPT [56] heads are employed to predict the forward and backward motions, respectively. Here, we define the forward/backward velocities $\{\boldsymbol{v}_{i}^{+},\boldsymbol{v}_{i}^{-}\}$ as the 3D displacements from the current frame to the next/previous frame in the camera coordinate.
@@
The multiple encoders for multi-modal conditions are implemented with 1) VAE [65] encoder for RGB images and depth maps, 2) convolutional layers with $8\times$ spatial and $4\times$ temporal compression ratio for masks and plüker embeddings. During the generation training stage, only convolutional layers are trainable while the VAE encoder is frozen.
@@ Appendix B Training Details
To ensure compatibility with the patch size of DINOv2 [54] in the reconstruction model ($\times 14$ downsampling) and the VAE in the generation model ($\times 8$ compression), we resize all input videos to have a longest edge of 560 pixels during reconstruction training, and a fixed resolution of $336\times 560$ during generation training.
@@
PointOdyssey [93] 131
DynamicReplica [34] 483
Kubric [21] 5.7K
Spring [53] 37
VKITTI2 [8] 50
Waymo [63] 798
TartanAir [72] 369
BEDLAM [6] 10.4K
MVS-Synth [29] 120
GFIE [27] 81
HOI4D [48] 3.0K
CoP3D [61] 2.8K
DL3DV [44] 6.4K
Scannet++ [85] 853
ARKitScenes [4] 4.5K
HyperSim [59] 457
MapFree [1] 460
SpatialVID [66] 371.3K
@@ Appendix D Evaluation Protocol
We evaluate static reconstruction performance on VRNeRF [77] and Scannet++ [85].
@@
For dynamic reconstruction on ADT [55], we follow 4DGT [78] to evaluate the same 4 scenes:
@@
For DyCheck [20], we evaluate 5 scenes (apple, block, paper-windmill, spin, teddy). We sample 64 consecutive timestamps for each scene, using 32 frames (stride 2) from a casually-captured video (camera 0) for reconstruction and the complete 64 frames from another fixed-camera video (camera 1) for testing.
[…]
AnyV2V [38] 0.071 24.89 5.04 6.11
Wan-Edit [39] 0.013 26.39 6.54 3.07
VACE [33] 0.015 26.92 4.37 4.30
Ours 0.018 26.66 5.13 0.49
Table S2: Video editing evaluation on FiVE [39].
SpatialTracker [75] St4RTrack [19] Ours
[…]
Table S3: 3D tracking evaluation on DriveTrack through TAPVid-3D [37]. The prediction of SpatialTracker is offered by TAPVid-3D and St4RTrack is predicted from its official codebase.
Appendix E Downstream Task Evaluation
In the main paper, we qualitatively demonstrate several downstream applications of NeoVerse. Here, we provide quantitative evaluations on two representative tasks: video editing and 3D tracking.
Video editing.
We evaluate video editing on the FiVE [39] benchmark and compare with AnyV2V [38], Wan-Edit [39], and VACE [33]. As shown in Tab. S2, although NeoVerse is not specifically designed for video editing, it achieves competitive performance while being significantly faster (0.49 seconds per frame vs. 3.07–6.11 seconds per frame for other methods).
3D tracking.
We evaluate 3D tracking on the DriveTrack subset of TAPVid-3D [37] and compare with SpatialTracker [75] and St4RTrack [19]. As shown in Tab. S3, it demonstrates that the 3D flow predicted by our reconstruction model provides reliable 3D correspondences.
Appendix F Discussion on Linear Motion Assumption
As described in Eq. 3, our method assumes approximately linear motion between adjacent key frames for Gaussian interpolation. While this is a simplified assumption, it does not negatively affect the generation quality for the following reasons. During training, we reconstruct from sparse key frames and render to all frames. The less-accurate non-keyframe renderings naturally serve as a form of temporal degradation, encouraging the generation model to learn to produce high-quality videos with non-linear motions from degraded renderings. During inference, users can input all frames to ensure reliable renderings when needed. Moreover, linear motion is a common and reasonable assumption adopted by prior works such as 4DGT [78], as real-world motion within short intervals between adjacent frames is generally well-approximated by linear interpolation.
Appendix G Limitations and Failure Cases
@@
Appendix H Additional Qualitative Results